### v1

@@ Abstract
Results: In simulations, CNN-based methods achieved up to 19% improvement in SSIM and 13% reduction in MAE compared to RL-PRC. The Two-Channel model showed the highest CR and CNR values, recovering lung activity with 97% agreement to the ground truth, compared to 77% with RL-PRC. Noise remained stable for CNN models ( 5.9%), whereas RL-PRC increased noise by 5.8%. In preclinical acquisitions, the Two-Channel model achieved the highest CNR in different tissues, while maintaining the lowest noise level (9.6%). Although no ground truth was available for real data, analysis confirmed superior tumor delineation and reduced spillover artifacts with the Two-Channel model.
@@ 1 Background
The use of 68Ga in PET imaging is growing significantly due to its versatile applications in both preclinical and clinical applications [1], the availability of 68Ge/68Ga generators [2], and the approval of 68Ga-labeled tracers for clinical use [3]. This radionuclide is particularly advantageous for labeling both small compounds and macromolecules [4], enabling a wide range of diagnostic uses, including cancer imaging with 68Ga-DOTATATE for neuroendocrine tumors [5, 6] and 68Ga-PSMA for prostate-specific membrane antigen (PSMA) targeting in prostate cancer [7, 8, 9]. Additionally, 68Ga is employed for infection and inflammation imaging [10], cardiac imaging for myocardial perfusion [11], and bone imaging for detecting metastases in prostate cancer [12]. However, 68Ga emits positrons with an average large energy compared to other standard PET radionuclides such as 18F and 11C (see Table 1), resulting in one of the main limiting factors of spatial resolution and accurate quantification in PET [13]. Accordingly, authors like L.M. Carter et al. [14] have characterized the positron range (PR) of standard and non-standard PET radionuclides to better assess their impact on image quality, which becomes more important as the intrinsic resolution of the state-of-the-art PET scanners improves.
Table 1: $\beta^{+}$ Yield, Energy, and Positron Range for 18F, 11C and 68Ga [15].
@@
PR is defined as the distance traveled by the positrons from the emission to the annihilation points, results in anisotropic blurring in reconstructed PET images. PR is influenced by both the kinetic energy spectrum of the emitted positrons (which is isotope-dependent) and the electronic density of the tissue through which the positron travels before annihilation. Higher positron kinetic energies and lower tissue densities lead to larger PR [16]. The mean energy of the positrons emitted by 68Ga is 836.0 keV, significantly higher than that of the commonly used 18F (249.3 keV) [15]. This energy difference results in a substantially increased PR for 68Ga, particularly evident in low-density tissues such as the lungs. As shown in Table 1, the mean PR in lung tissue increases from 1.85 mm for 18F to 8.86 mm for 68Ga, with the maximum PR expanding from 7.49 mm to 27.1 mm, respectively. This trend is also present in other tissues, such as water, which shows an increase from 0.57 mm to 2.69 mm (mean PR) and bone, from 0.32 mm to 1.44 mm. The pronounced effect highlights the need for positron range correction (PRC) when using high-energy positron emitters like 68Ga. PRC is essential to produce quantitative PET images with spatial resolution comparable to those obtained with 18F or 11C, ensuring accurate localization and quantification of radiotracer uptake, especially in heterogeneous environments.
@@
Several approaches have been proposed to correct the blurring caused by PR within the reconstruction process [17]. Derenzo [18] proposed reducing PR effect in the projection domain by Fourier deconvolution combined with the Filtered Back Projection (FBP) reconstruction. However, this approach does not consider the heterogeneous nature of PR in different materials and increases noise through deconvolution, especially in narrow projection bins and for isotopes with high positron average energies [18]. Some PRC methods are based on modeling the Point Spread Function (PSF) within the System Response Matrix (SRM) used in iterative reconstruction [19, 20, 21]. While most reconstruction algorithms account for the PR of 18F in water by modeling a default PSF of the system, modeling the PR of non-standard PET isotopes, such as 68Ga, remains challenging because the PSF or SRM must be modified. Additionally, as PR depends on the specific local tissue properties, realistic PR models require the SRM to be evaluated at each acquisition, which is not computationally efficient [22]. Other methods use material-dependent isotropic filters derived from CT or MRI anatomical images to deconvolve the PR effect with a combination of kernels [23]. Even these methods have shown promising results, they still present limitations especially in the boundaries of high density regions such as bone tissue [17].
@@
Recently, Gavriilidis et al. proposed a tissue-independent 68Ga-specific PRC using a convolution kernel modeling the PR of 68Ga in water using Monte Carlo simulations [24, 25]. This kernel is integrated with the OSEM and Q.Clear iterative reconstruction algorithms, incorporating point spread function (PSF) modeling and time-of-flight (ToF) information. Despite its potential, tissue-independent PRC can lead to inaccuracies, especially at tissue boundaries, due to over- or under-correction in different tissue types [24].
@@
Vass and Reader proposed a method to address spatial resolution losses in reconstructed images caused by effects like PR [26]. This method involves forward projecting the reconstructed image using a user-defined virtual scanner geometry to generate synthetic sinogram data. The synthetic sinogram data are then used as input for image reconstruction, integrating a modeled PSF to correct PR-induced resolution degradation.
@@
PRC can be also dealt with as a post-processing step by deconvolution of the reconstructed images. A simple and fast strategy is to simulate a kernel that characterizes the blur due to PR in a specific tissues (e.g. water, soft tissue, bone, lung, etc.) and perform a Richardson-Lucy deconvolution using the kernel to de-blur the reconstructed image [27]. However, this method recovers the loss of spatial resolution without considering the anisotropic nature of PR in the boundaries of tissues and often increases image noise caused by the deconvolution process [17].
@@
With the popularization of artificial intelligence, methods using convolutional neural networks (CNNs) have been proposed to reduce the PR effect when working with 68Ga. Yang [28] tested three different CNN models for PRC in preclinical PET imaging of 68Ga. The study used a preclinical PET scanner and 30 phantom models in Monte Carlo simulations, running each model twice: once with 68Ga (for CNN input images) and once with back-to-back 511-keV gamma rays. Although the recovery coefficient and spill-over ratio improved after correction, the evaluation was only done on phantom simulations and did not include information from density in the training, which may affect if it is applied to real acquisition. In previous works [22], Herraiz et al. demonstrated the potential of employing a U-Net architecture [29] trained on simulated data to achieve fast and accurate PRC, incorporating anatomical information from the $\mu$-map during training. However, while the results showed improved performance when this information was included, no specific constraint was applied to ensure the proper use of the anatomical data.
@@
In this work, we build on previous work [22] including a novel network working on 3D patches instead of 2D slices and a tissue-dependent loss function to enforce consistency with the anatomic information provided by the $\mu$-map. We also significantly expanded the training dataset, improving the model’s ability to generalize across diverse scenarios. The use of multi-scale volumetric patches better describe the localized three-dimensional effect of PR when compared to 2D slices approaches. For the modified loss function, we explored three distinct strategies for integrating anatomical information from the $\mu$-map: indirectly (Single-Channel), as an additional input channel (Two-Channel), and through a dedicated encoder (DualEncoder). To benchmark performance, we compared our method with the standard Richardson-Lucy post-processing algorithm [27]. Notably, we also evaluated the accuracy and impact of our approach using 68Ga-FH and 68Ga-PSMA mice acquisitions, showcasing the impact of PRC on real studies.
@@ 2.1 Synthetic PET data
Computational animal models are commonly used to perform simulations with realistic materials and anatomies of animals in radiation transport and radiotracer biodistribution studies. In this work, MOBY [30] and Digimouse [31] whole-body numerical phantoms were employed to simulate activity, material, and density distributions. The original Digimouse model was modified to introduce a variety of tumor shapes and tissues, such as the myocardium, prostate, and thyroid, which were not included in the original implementation. Differently to our previous work [22], which relied on PET/CT images from a single database, the use of different atlas provides increased control over tissue characterization and allows us to generate a variety of activity distributions within the body based on radiotracers used in PET imaging (68Ga-PSMA, 68Ga-DOTATOC, 68Ga-DOTATATE, 18F-FDG), as well as additional synthetic distributions.
@@
To simulate the PR effects, an adapted PenEasy [32] version was developed to generate positron annihilation distributions based on the initial positron emission distribution, using material and density maps from MOBY and Digimouse models. PenEasy tracks the path of each positron until its annihilation, taking into account their energy and all materials within the path using positron energy distributions simulated with PenNuc [33], which considers all potential decay branches and nuclear properties for a wide range of radionuclides. The formation of positronium was not considered and positrons were assumed to annihilate in the voxel where their energy dropped below 100 eV.
@@
The positron annihilation distributions from the adapted PenEasy were introduced into MCGPU-PET [34] to produce fast and realistic simulations based on the geometry of the Inveon PET/CT scanner [35]. MCGPU-PET outputs were stored in sinograms with 147×168×1293 bins, maximum ring difference of 79, an axial compression factor (span) of 11, and a radial bin size of 0.795 mm. These sinograms were reconstructed using GFIRST [36], an PSF-3D-OSEM algorithm implemented in CUDA using 60 iterations and 1 subset, to obtain 154×154×159 (x,y,z voxels) reconstructed images with a voxel size of 0.776×0.776×0.795 mm. Each reconstructed image was based on simulations of 2 million events, ensuring a realistic representation of PET acquisition conditions.
@@
The attenuation were obtained from MOBY and Digimouse phantoms and converted to mm-1 by defining three different materials and their respective reference densities [37]: air (CT = 0, $\rho$ = 0.0012 g/cm3), soft tissue (CT $\sim$ 110, $\rho$ = 1.0 g/cm3), and bone (CT = 255, $\rho$ $\sim$ 1.85 g/cm3). A piecewise linear mapping was applied: one segment interpolating from air to water, and another from water to bone, yielding a continuous density map across the CT volume. This map was then converted into a $\mu$-map by scaling the densities with linear attenuation coefficient for 511 keV gammas.
@@
The workflow was implemented for both 18F and 68Ga, resulting in a database that includes two reconstructed PET images and the $\mu$-map image for each numerical phantom. We assume that the 18F PRC correction is included in the PSF of the system, which also includes other effects such as non-collinearity.
@@ 2.2 CNNs models
Three 3D residual encoder-decoder convolutional neural networks (3D RED-CNN) were developed. The model architectures are based on fully convolutional 2D CNN with shortcut connections between the encoder and the decoder proposed by Chen et al. [38], which has been demonstrated to outperform more complex models in low-dose CT denoising [39].
Two key modifications were introduced to the original design: (a) 3D convolutional and deconvolutional layers replaced 2D layers to leverage 3D spatial information and mitigate artifacts between adjacent 2D slices [40]; (b) the incorporation of anatomical information to enhance the network’s performance. The integration of $\mu$-map information was evaluated using three different architectures, as shown in Figure 1. In the first NN architecture, the Single-Channel 3D RED-CNN processes only PET input, while the $\mu$-map information is indirectly incorporated by influencing the loss function during training (Fig. 1A). In the second NN, the Two-Channel 3D RED-CNN builds upon the Single-channel architecture by adding an additional input channel, allowing direct incorporation of anatomical information from the $\mu$-map alongside the PET input (Fig. 1B). In the third architecture, the DualEncoder 3D RED-CNN processes the $\mu$-map information through an additional encoder pathway, enabling a deeper integration of anatomical features into the network’s feature extraction process (Fig. 1C). These variations aim to explore the impact of different methods of incorporating anatomical information on the network’s performance.
The Single-channel model, composed of convolutional layers and ReLU activation functions, is inherently scale-independent due to the shift-invariant nature of convolutions [41] and the positive homogeneity of ReLU, requiring no explicit normalization during inference. In contrast, the Two-channel and DualEncoder models combine PET and $\mu$-map inputs, which differ in dynamic range and intensity distribution, making normalization essential. To ensure consistency, each PET image was normalized by the maximum of the 18F image after removing outliers, while $\mu$-map images were normalized by their own maximum values.
@@ 2.3 Tissue-dependent loss function
We use a loss function consisting of two components: (a) the Mean Absolute Error (MAE), a voxel-wise loss quantifying the error between the predicted image generated by the model and the ground truth PET image, and (b) a regularization term incorporating $\mu$-map information. To achieve this, Mutual Information (MI), a measure of the shared information between two random variables $X$ and $Y$ [42], can be used to force the NN to use the anatomical information. Formally, MI is defined as follows [43],
$\mathrm{MI}(X;Y)=\sum_{x\in X}\sum_{y\in Y}P(x,y)\cdot log\left(\frac{P(x,y)}{P(x)\cdot P(y)}\right)$ (1)
where $P(x,y)$ is the joint probability distribution of $X$ and $Y$, and $P(x)\cdot P(y)$ is the product of the marginal distributions. When dealing with images, $X$ and $Y$ are sets of pixels $x$ and $y$, respectively. $\mathrm{MI}$ is commonly used in multimodal registration for PET/CT images [44].
Based on this concept, the differentiable global mutual information loss proposed by the open-source MONAI framework for DL in healthcare [45] was selected, which allows the implementation of MI as a loss function,

### v2

@@ Abstract
Results: In simulations, CNN-based methods achieved up to 19% improvement in SSIM and 13% reduction in MAE compared to RL-PRC. The Two-Channel model showed the highest CR and CNR values, recovering lung activity with 97% agreement to the ground truth, compared to 77% with RL-PRC. Noise remained stable for CNN models ($\sim$5.9%), whereas RL-PRC increased noise by 5.8%. In preclinical acquisitions, the Two-Channel model achieved the highest CNR in different tissues, while maintaining the lowest noise level (9.6%). Although no ground truth was available for real data, analysis confirmed superior tumor delineation and reduced spillover artifacts with the Two-Channel model.
@@ 1 Background
The use of 68Ga in PET imaging is growing significantly due to its versatile applications in preclinical and clinical studies [baum2012theranostics], the availability of 68Ge/68Ga generators [roesch2012maturation], and the approval of 68Ga-labeled tracers for clinical use [hennrich202168ga]. This radionuclide is particularly advantageous for labeling both small compounds and macromolecules [conti2016physics], enabling a wide range of diagnostic uses, including cancer imaging with 68Ga-DOTATATE for neuroendocrine tumors [mojtahedi2014value, haug2012role] and 68Ga-PSMA for prostate-specific membrane antigen (PSMA) targeting in prostate cancer [bois202068ga, lenzo2018review, han2018impact]. Additionally, 68Ga is employed for infection and inflammation imaging [xu2020research], cardiac imaging for myocardial perfusion [autio202068], and bone imaging for detecting metastases in prostate cancer [sachpekidis201868]. However, 68Ga emits positrons with an average large energy compared to other standard PET radionuclides such as 18F and 11C (see Table 1), resulting in one of the main limiting factors of spatial resolution and accurate quantification in PET [levin1999calculation]. Accordingly, authors like L.M. Carter et al. [carter2020impact] have characterized the positron range (PR) of standard and non-standard PET radionuclides to better assess their impact on image quality, which becomes more important as the intrinsic resolution of the state-of-the-art PET scanners improves.
Table 1: $\beta^{+}$ Yield, Energy, and Positron Range for 18F, 11C and 68Ga [gonzalez2014positron].
@@
PR is defined as the distance traveled by the positrons from the emission to the annihilation points, resulting in anisotropic blurring in reconstructed PET images. PR is influenced by both the kinetic energy spectrum of the emitted positrons (which is isotope-dependent) and the electronic density of the tissue through which the positron travels before annihilation. Higher positron kinetic energies and lower tissue densities lead to larger PR [cal2013positron]. The mean energy of the positrons emitted by 68Ga is 836.0 keV, significantly higher than that of the commonly used 18F (249.3 keV) [gonzalez2014positron]. This energy difference results in a substantially increased PR for 68Ga, particularly evident in low-density tissues such as the lungs. As shown in Table 1, the mean PR in lung tissue increases from 1.85 mm for 18F to 8.86 mm for 68Ga, with the maximum PR expanding from 7.49 mm to 27.1 mm, respectively. This trend is also present in other tissues, such as water, which shows an increase from 0.57 mm to 2.69 mm (mean PR) and bone, from 0.32 mm to 1.44 mm. The pronounced effect highlights the need for positron range correction (PRC) when using high-energy positron emitters like 68Ga. PRC is essential to produce quantitative PET images with spatial resolution comparable to those obtained with 18F or 11C, ensuring accurate localization and quantification of radiotracer uptake, especially in heterogeneous environments.
@@
Several approaches have been proposed to correct the blurring caused by PR within the reconstruction process [gavriilidis2022positron]. Derenzo [derenzo1986mathematical] proposed reducing PR effect in the projection domain by Fourier deconvolution combined with the Filtered Back Projection (FBP) reconstruction. However, this approach does not consider the heterogeneous nature of PR in different materials and increases noise through deconvolution, especially in narrow projection bins and for isotopes with high positron average energies [derenzo1986mathematical]. Some PRC methods are based on modeling the Point Spread Function (PSF) within the System Response Matrix (SRM) used in iterative reconstruction [cal2015tissue, cal2018improving, kertesz2022implementation]. While most reconstruction algorithms account for the PR of 18F in water by modeling a default PSF of the system, modeling the PR of non-standard PET isotopes, such as 68Ga, remains challenging because the PSF or SRM must be modified. Additionally, since PR depends on the specific local tissue properties, realistic PR models require the SRM to be evaluated at each acquisition, which is not computationally efficient [lopez2020deep]. Other methods use material-dependent isotropic filters derived from CT or MRI anatomical images to deconvolve the PR effect with a combination of kernels [cal2011study]. Even though these methods have shown promising results, they still present limitations, especially at the boundaries of high-density regions such as bone tissue [gavriilidis2022positron].
@@
Recently, Gavriilidis et al. proposed a tissue-independent 68Ga-specific PRC using a convolution kernel modeling the PR of 68Ga in water using Monte Carlo simulations [gavriilidis2024performance, gavriilidis2025impact]. This kernel is integrated with the OSEM and Q.Clear iterative reconstruction algorithms, incorporating point spread function (PSF) modeling and time-of-flight (ToF) information. Despite its potential, tissue-independent PRC can lead to inaccuracies, especially at tissue boundaries, due to over- or under-correction in different tissue types [gavriilidis2024performance].
@@
Vass and Reader proposed a method to address spatial resolution losses in reconstructed images caused by effects like PR [vass2023synthesized]. This method involves forward projecting the reconstructed image using a user-defined virtual scanner geometry to generate synthetic sinogram data. The synthetic sinogram data are then used as input for image reconstruction, integrating a modeled PSF to correct PR-induced resolution degradation.
@@
PRC can also be dealt with as a post-processing step by deconvolution of the reconstructed images. A simple and fast strategy is to simulate a kernel that characterizes the blur due to PR in a specific tissues (e.g. water, soft tissue, bone, lung, etc.) and perform a Richardson-Lucy deconvolution using the kernel to de-blur the reconstructed image [rukiah2018investigation]. However, this method recovers the loss of spatial resolution without considering the anisotropic nature of PR in the boundaries of tissues and often increases image noise caused by the deconvolution process [gavriilidis2022positron].
@@
With the popularization of artificial intelligence, methods using convolutional neural networks (CNNs) have been proposed to reduce the PR effect when working with 68Ga. Yang [yang2021compensating] tested three different CNN models for PRC in preclinical PET imaging of 68Ga. The study used a preclinical PET scanner and 30 phantom models in Monte Carlo simulations, running each model twice: once with 68Ga (for CNN input images) and once with back-to-back 511-keV gamma rays. Although the recovery coefficient and spill-over ratio improved after correction, the evaluation was only done on phantom simulations and did not include information from density in the training, which may affect if it is applied to real acquisition. In previous works [lopez2020deep], Herraiz et al. demonstrated the potential of employing a U-Net architecture [ronneberger2015u] trained on simulated data to achieve fast and accurate PRC, incorporating anatomical information from the $\mu$-map during training. However, while the results showed improved performance when this information was included, no specific constraint was applied to ensure the proper use of the anatomical data.
@@
In this work, we build on previous work [lopez2020deep] including a novel network working on 3D patches instead of 2D slices and a tissue-dependent loss function to enforce consistency with the anatomic information provided by the $\mu$-map. We also significantly expanded the training dataset, improving the model’s ability to generalize across diverse scenarios. The use of multi-scale volumetric patches better describe the localized three-dimensional effect of PR when compared to 2D slices approaches. For the modified loss function, we explored three distinct strategies for integrating anatomical information from the $\mu$-map: indirectly (Single-Channel), as an additional input channel (Two-Channel), and through a dedicated encoder (DualEncoder). To benchmark performance, we compared our method with the standard Richardson-Lucy post-processing algorithm [rukiah2018investigation]. Notably, we also evaluated the accuracy and impact of our approach using 68Ga-FH and 68Ga-PSMA-617 mice acquisitions, showcasing the impact of PRC on real studies.
@@ 2.1 Synthetic PET data
Computational animal models are commonly used to perform simulations with realistic materials and anatomies of animals in radiation transport and radiotracer biodistribution studies. In this work, MOBY [segars2004development] and Digimouse [dogdas2007digimouse] whole-body numerical phantoms were employed to simulate activity, material, and density distributions. The original Digimouse model was modified to introduce a variety of tumor shapes and tissues, such as the myocardium, prostate, and thyroid, which were not included in the original implementation. Unlike our previous work [lopez2020deep], which relied on PET/CT images from a single database, the use of different atlas provides increased control over tissue characterization and allows us to generate a variety of activity distributions within the body based on radiotracers used in PET imaging (68Ga-PSMA, 68Ga-DOTATOC, 68Ga-DOTATATE, 18F-FDG), as well as additional synthetic distributions.
@@
To simulate the PR effects, an adapted PenEasy [sempau2006configuration] version was developed to generate positron annihilation distributions based on the initial positron emission distribution, using material and density maps from MOBY and Digimouse models. PenEasy tracks the path of each positron until its annihilation, taking into account their energy and all materials within the path using positron energy distributions simulated with PenNuc [garcia2019pennuc], which considers all potential decay branches and nuclear properties for a wide range of radionuclides. The formation of positronium was not considered and positrons were assumed to annihilate in the voxel where their energy dropped below 100 eV.
@@
The positron annihilation distributions from the adapted PenEasy were introduced into MCGPU-PET [herraiz2024mcgpu] to produce fast and realistic simulations based on the geometry of the Inveon PET/CT scanner [constantinescu2009performance]. MCGPU-PET outputs were stored in sinograms with 147×168×1293 bins, maximum ring difference of 79, an axial compression factor (span) of 11, and a radial bin size of 0.795 mm. These sinograms were reconstructed using GFIRST [herraiz2011gpu], an PSF-3D-OSEM algorithm implemented in CUDA using 60 iterations and 1 subset, to obtain 154×154×159 (x,y,z voxels) reconstructed images with a voxel size of 0.776×0.776×0.795 mm. Each reconstructed image was based on simulations of 2 million events, ensuring a realistic representation of PET acquisition conditions.
@@
The attenuation maps were obtained from MOBY and Digimouse phantoms and converted to mm-1 by defining three different materials and their respective reference densities [berger_stopping-power_2017]: air (CT = 0, $\rho$ = 0.0012 g/cm3), soft tissue (CT $\sim$110, $\rho$ = 1.0 g/cm3), and bone (CT = 255, $\rho$ $\sim$1.85 g/cm3). A piecewise linear mapping was applied: one segment interpolating from air to water, and another from water to bone, yielding a continuous density map across the CT volume. This map was then converted into a $\mu$-map by scaling the densities with linear attenuation coefficient for 511 keV gammas.
@@
The workflow was implemented for both 18F and 68Ga, resulting in a database that includes two reconstructed PET images and the $\mu$-map image for each numerical phantom. We assume that the 18F PRC correction is included in the PSF of the system, which also includes other effects such as non-collinearity. Representative examples of the simulated, synthetic, and real PET datasets used in this study are shown in supplementary material. To improve reproducibility, 15 Digimouse-based numerical phantoms generated for this work have been uploaded to a Zenodo repository [encina_baranda]. MOBY-based datasets are not included due to license restrictions, as well as real data acquisitions.
@@ 2.2 CNNs models
Three 3D residual encoder-decoder convolutional neural networks (3D RED-CNN) were developed. The model architectures are based on fully convolutional 2D CNN with shortcut connections between the encoder and the decoder proposed by Chen et al. [chen2017low], which has been demonstrated to outperform more complex models in low-dose CT denoising [eulig2024benchmarking].
Two key modifications were introduced to the original design: (a) 3D convolutional and deconvolutional layers replaced 2D layers to leverage 3D spatial information and mitigate artifacts between adjacent 2D slices [chen2023total]; (b) the incorporation of anatomical information to enhance the network’s performance. The integration of $\mu$-map information was evaluated using three different architectures, as shown in Figure 1. In the first NN architecture, the Single-Channel 3D RED-CNN processes only PET input, while the $\mu$-map information is indirectly incorporated by influencing the loss function during training (Fig. 1A). In the second NN, the Two-Channel 3D RED-CNN builds upon the Single-Channel architecture by adding an additional input channel, allowing direct incorporation of anatomical information from the $\mu$-map alongside the PET input (Fig. 1B). In the third architecture, the DualEncoder 3D RED-CNN processes the $\mu$-map information through an additional encoder pathway, enabling a deeper integration of anatomical features into the network’s feature extraction process (Fig. 1C). These variations aim to explore the impact of different methods of incorporating anatomical information on the network’s performance.
@@ 2.3 Tissue-dependent loss function
We use a loss function consisting of two components: (a) the Mean Absolute Error (MAE), a voxel-wise loss quantifying the error between the predicted image generated by the model and the ground truth PET image, and (b) a regularization term incorporating $\mu$-map information. To achieve this, Mutual Information (MI), a measure of the shared information between two random variables $X$ and $Y$ [cover1999elements], can be used to force the NN to use the anatomical information. Formally, MI is defined as follows [vergara2014review],
$\mathrm{MI}(X;Y)=\sum_{x\in X}\sum_{y\in Y}P(x,y)\cdot log\left(\frac{P(x,y)}{P(x)\cdot P(y)}\right)\$(1)
where $P(x,y)$ is the joint probability distribution of $X$ and $Y$, and $P(x)\cdot P(y)$ is the product of the marginal distributions. When dealing with images, $X$ and $Y$ are sets of pixels $x$ and $y$, respectively. $\mathrm{MI}$ is commonly used in multimodal registration for PET/CT images [maes1997multimodality].
Based on this concept, the differentiable global mutual information loss proposed by the open-source MONAI framework for DL in healthcare [cardoso2022monai] was selected, which allows the implementation of MI as a loss function,